OpenStudy (anonymous):

solve the equation for F in terms of Y C=5/9(f-32)

OpenStudy (one098):

What's y supposed to be?

OpenStudy (anonymous):

sorry i ment F in terms of C

OpenStudy (one098):

Okay, you need the value for f.. Do they give the temperature in F?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the topic is supposed to be rewriting equations

OpenStudy (one098):

Okay you multiply both sides by nine, to clear out the denominator. Then divide both sides by 5. Then add the 32 to both sides.

OpenStudy (anonymous):

but arent u supposed to do the distributive property since f-32 are in parentheses

OpenStudy (one098):

No because you are isolating the f and turning the equation around and moving everything to the other side..

OpenStudy (anonymous):

can you type your work so i can understand better

OpenStudy (one098):

\[C=\frac{ 5 }{ 9 }(f-32)\] \[9\times C=\frac{ 5 }{ 9 }(f-32)\times9\] \[9\times C=5\times(F-32)\] \[\frac{ 9C }{ 5 }=\frac{ 5\times(F-32) }{ 5 }\] \[\frac{ 9 }{ 5 } \times C=F-32\] \[\frac{ 9 }{ 5 } \times C + 32=F-32+32\] \[F=\frac{ 9 }{ 5 } \times C=32\]